OpenStudy (anonymous):

help please???? 4. Determine two different values of “b” in x2 + bx + 30 so that the expression can be factored into the product of two binomials. Explain how you determined those values and show each factorization. Explain how your process would change if the expression was 2x2 + bx + 30.

OpenStudy (anonymous):

a*c=1*30=30 b,can be 5+6,3+10,2+15,1+30 5*6=30 3*10=30 2*15=30 1*30=30 i have told you four values for second part 2*30=60 b can be 1+60 2+30 3+20 4+15 5+12 6+10